For a 1cm pathlength, the absorbance at 260nm (A260) equals 1.0 for a 50 microgram/mL solution of double stranded DNA. What is (i) the concentration and (ii) the amount of nucleic acid in 750 microlitre of a solution of double stranded DNA with an A260 of 0.46?
Answer #1

For a 1cm pathlength, the absorbance at 260nm (A260) equals 1.0 for a 50 microgram/mL solution of double stranded DNA.

we have : A = e*C*l

e = molar absorption coefficient for double stranded DNA ( at 260 nm)

C = concentration = 50 ug/mL    ( microgram /mL)

l = path length = 1.0 cm

A = absorbance = 1.0

we have "

e = A /C*l = 0.02 (ug/mL)-1 cm-1

A sample solution of double stranded DNA with an A260 : 0.46

(i) the concentration :

A = e*C*l = 0.46

0.02 (ug/mL)-1 cm-1 * C *1.0 cm = 0.46

C = 23 ug/mL

(ii) the amount of nucleic acid in 750 microlitre of sample solution :

1000 microlitre = 1 mL

thus., 750 microlitre = 0.75 mL

thus , amount of nucleic acid in 750 microlitre of sample solution = 0.75 mL*23 ug/mL

                      = 17.25 microgram